# Loc2: Interpretable Cross-View Localization via Depth-Lifted Local Feature Matching

### v2

@@ Abstract
We propose an accurate and interpretable fine-grained cross-view localization method that estimates the 3 Degrees of Freedom (DoF) pose of a ground-level image by matching its local features with a reference aerial image. Unlike prior approaches that rely on global descriptors or bird’s-eye-view (BEV) transformations, our method directly learns ground–aerial image-plane correspondences using weak supervision from camera poses. The matched ground points are lifted into BEV space with monocular depth predictions, and scale-aware Procrustes alignment is then applied to estimate camera rotation, translation, and optionally the scale between relative depth and the aerial metric space. This formulation is lightweight, end-to-end trainable, and requires no pixel-level annotations. Experiments show state-of-the-art accuracy in challenging scenarios such as cross-area testing and unknown orientation. Furthermore, our method offers strong interpretability: correspondence quality directly reflects localization accuracy and enables outlier rejection via RANSAC, while overlaying the re-scaled ground layout on the aerial image provides an intuitive visual cue of localization accuracy.
@@ 1 Introduction
Recently, FG2 (Xia & Alahi, 2025) demonstrated the feasibility of establishing local feature correspondences between ground and aerial images by propagating BEV point matches to the image plane. However, performing the matching in BEV is suboptimal, since warping a ground image into BEV inevitably introduces ray-directional distortions (Song et al., 2024) and discards information along the height dimension (Xia & Alahi, 2025; Wang et al., 2024a). This loss of information hinders matching to aerial images and, as a result, degrades localization performance, especially in challenging settings such as unknown camera orientation.
Therefore, we propose to establish correspondences directly between ground and aerial images. Instead of relying on pixel-level annotations (Leroy et al., 2024; Wang et al., 2025), our approach is weakly supervised using only 3-DoF camera poses. To infer the pose from image-plane correspondences, we lift the sampled ground points into the BEV space with the help of monocular depth models (Yang et al., 2024a; b; Piccinelli et al., 2025; Wang & Liu, 2024). In unseen environments, usually only the non-metric relative depth prediction, i.e., depth up to an unknown scale, is reliable. Therefore, our method supports both metric and relative depth, and jointly estimates the camera pose and depth scale using scale-aware Procrustes alignment (Umeyama, 1991).
@@ 2 Related Work
Fine-grained cross-view localization is challenging due to the drastic viewpoint differences and asynchronous capture times between ground and aerial imagery. State-of-the-art approaches tackle this domain gap primarily by learning a deep network for camera pose estimation. Global descriptor-based methods (Xia et al., 2022; Lentsch et al., 2023; Xia et al., 2023) leverage contrastive learning to pull the ground and aerial features to a common representation space, and compare the ground image descriptor to aerial descriptors extracted at different candidate poses. Geometry transformation-based methods (Fervers et al., 2023b; Wang et al., 2023; Sarlin et al., 2024; Wang et al., 2024a; Song et al., 2024; Shi et al., 2023; Wang et al., 2024b; Fervers et al., 2023a; Xia & Alahi, 2025) first warp the ground image into a BEV representation, and then either perform a sliding-window search over the aerial image to find the best alignment (Sarlin et al., 2024; Fervers et al., 2023b; Shi et al., 2023) or estimate the relative pose by matching features between the BEV and aerial view (Song et al., 2024; Wang et al., 2024b; 2023; a; Xia & Alahi, 2025). Despite the increasing localization accuracy, most methods do not explicitly identify matched local features across views, offering limited interpretability. A recent work (Xia & Alahi, 2025) demonstrated, for the first time, the ability to find ground-aerial local feature correspondences by defining a dense 3D point cloud and querying image features for each point. However, this process is inefficient due to the sparse nature of 3D space, and this method performs poorly when camera orientation is unknown.
@@ 3.2 Camera Pose Estimation
Coordinate assignment: Given the $N$ sampled correspondences, we leverage off-the-shelf monocular depth estimators to assign coordinates to these correspondences, which are then used to estimate the camera pose. Importantly, monocular depth prediction is a geometrically ill-posed problem, as there is no unique mapping from a single image to absolute metric depth. State-of-the-art methods either learn generalizable metric depth from diverse training data (Piccinelli et al., 2025; Zhu et al., 2024), or predict depth that is accurate only up to an unknown per-image scale (Wang & Liu, 2024; Yang et al., 2024a). To generalize across diverse scenarios and image types, our method accommodates both metric and relative depth, and explicitly estimates a scale factor $s$ to convert relative depth into metric space  [^1] Recent work (Yu et al., 2025) models relative depth using a scale and a shift. As our goal is not metric depth prediction, we follow Eigen et al. (2014); Ding et al. (2024) and consider only the scale for simplicity.. This scale estimation leverages the metric information, i.e., meters per pixel, available in geo-referenced aerial imagery.
@@
Scale-aware Procrustes alignment: As shown in Fig. 2 right, once we have the point correspondences $\{(x_{n}^{A},y_{n}^{A}),(x_{n}^{G},y_{n}^{G},z_{n}^{G})/s\}$, and their matching probabilities $w_{n}$, we can compute the rotation $\mathbf{R}$, translation $\mathbf{t}$, and the scale $s$ between the ground points and the metric space aerial points analytically in a differentiable manner using 2D scale-aware Procrustes alignment (Umeyama, 1991)  [^2] In Xia & Alahi (2025); Barroso-Laguna et al. (2024), all coordinates are in metric space, so they use orthogonal Procrustes (Gower, 1975) and ignore scale..
For simplicity, we denote the planar coordinates of all ground points as $\mathbf{P}=\mathbf{P}^{*}/s$, assuming they differ from the metric coordinates, $\mathbf{P^{*}}=\{(x^{G}_{1},y^{G}_{1}),...,(x_{n}^{G},y_{n}^{G})\}$, by an unknown scale $s$. The coordinates of all aerial points are denoted as $\mathbf{Q}$. The objective is to estimate the scale $s$, rotation matrix $\mathbf{R}$ and metric translation $\mathbf{t}$ that satisfy the transformation $\mathbf{Q}=s(\mathbf{R}\cdot\mathbf{P})+{\mathbf{t}}$.
@@
Next, we perform Singular Value Decomposition (SVD) on $\mathbf{C}$, where $\mathbf{C}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$, and the optimal rotation can be recovered as $\mathbf{R}=\mathbf{V}\mathbf{U}^{\top}$, which derives the estimated yaw orientation $o$. The scale and translation from the ground points $\mathbf{P}$ to the aerial points $\mathbf{Q}$ is then computed as,
@@
Recall that $\mathbf{P}=\mathbf{P}^{*}/s$. Substituting this into Eq. 4, the estimated scale $s^{*}$ can be expressed using the (unscaled) metric coordinates $\mathbf{P}^{*}$, with $\mathbf{\Sigma}^{*}$ being the singular value matrix computed from the SVD between $\mathbf{Q}$ and $\mathbf{P}^{*}$:
@@ 3.3 Model Supervision
We follow Xia & Alahi (2025) to provide supervision on both the pose and correspondences. For camera pose, we adopt the Virtual Correspondence Error (VCE) loss. Specifically, we define a set of $N_{v}$ virtual points $\mathbf{P}_{v}$ in a 2D metric space, where the hyperparameter $l$ controls the spatial extent of the space. We apply both the ground-truth transformation, $\mathbf{R}_{\text{gt}},\mathbf{t}_{\text{gt}}$, and the estimated transformation, $\mathbf{R},\mathbf{t}$, to these virtual points. The loss $\mathcal{L}_{\text{VCE}}$ then minimizes the mean Euclidean distance between the corresponding transformed points,
@@ 4 Experiment
We first introduce the datasets and metrics, followed by implementation details. We then compare our method with previous state-of-the-art vision-based methods and showcase our inference using different monocular depth predictors. Next, we present qualitative results and our ablation study.
@@ 4.1 Datasets and Evaluation Metrics
KITTI (Geiger et al., 2013) provides forward-facing ground-level images with a limited field of view. Aerial images for KITTI were collected by Shi & Li (2022), who also split the dataset into Training, Test1, and Test2 subsets. Test1 includes images from the same region as the Training set, while Test2 contains images from a different unseen region. We refer to Test1 and Test2 as the same-area and cross-area test sets, respectively. We use the metric depth predictions from DepthAnythingV2 (Yang et al., 2024b) for our experiments on KITTI.
VIGOR (Zhu et al., 2021) is a widely used cross-view localization dataset, containing ground-level panoramas and aerial images from four U.S. cities. It also provides same-area and cross-area splits. In the same-area split, training and test images are collected from all four cities, while in the cross-area split, the training and test sets come from two different cities. During training, we adopt a recent metric depth model, Unik3D (Piccinelli et al., 2025), which has shown strong performance on outdoor panoramic images. At test time, we report results using both Unik3D and relative depth models (Wang & Liu, 2024; Jiang et al., 2021; Wang et al., 2022). Additionally, we include results in the Appendix where both training and testing are performed using only relative depth.
@@ 4.2 Implementation Details
On both datasets, we use AdamW (Loshchilov & Hutter, 2019) with a learning rate of $1\times 10^{-4}$. Training is conducted on a single H100 GPU with a batch size of 80 for VIGOR and 224 for KITTI. We follow the setting from Xia & Alahi (2025): the temperature parameter $\tau$ in matching score computation is set to $0.1$, the aerial feature map $\mathbf{F}_{A}$ is resized to generate $41\times 41$ aerial points, $\mathcal{L}_{\text{VCE}}$ uses $N_{v}=10\times 10$ points, with $l$ set to 5 m, and $N=1024$ correspondences are sampled for pose estimation. For $\mathcal{L}_{\text{S2G}}$, negatives are 1 m away in planar distance from the projected location. We use $\beta=1$ for VIGOR and $\beta=0.1$ for KITTI. RANSAC (Fischler & Bolles, 1981) is applied on VIGOR, but omitted on KITTI as it did not yield improvements in localization accuracy. When using metric depth, we apply a maximum depth threshold of 35 m for VIGOR and 40 m for KITTI, setting the matching score of points exceeding this threshold to zero. For relative depth predictors, we apply a fixed initial scale (identified by visually inspecting a few examples) to all relative depth maps and then use the threshold to filter out matches corresponding to sky and distant objects.
@@ 4.3 Quantitative Results
KITTI: We compare our method with both global descriptor-based methods, CCVPE (Xia et al., 2023) and SliceMatch (Lentsch et al., 2023), and geometry transformation-based methods, GGCVT (Shi et al., 2023), HC-Net (Wang et al., 2024b), DenseFlow (Song et al., 2024), and FG2 (Xia & Alahi, 2025). As shown in Tab. 1, our method sets a new state-of-the-art in mean and median localization errors on the cross-area test under both $\pm 10^{\circ}$ and $\pm 180^{\circ}$ orientation noise. On the same-area test, although our method has higher error under $\pm 10^{\circ}$ noise, it shows a large improvement over the previous state-of-the-art method in the more challenging $\pm 180^{\circ}$ setting, reducing the mean error from 6.88 m to 1.85 m. For orientation prediction, our method is less accurate than the state of the art. Since our method computes orientation analytically from local feature correspondences, it cannot take full advantage of the prior that most ground images in KITTI are aligned with the road direction. Under $\pm 10^{\circ}$ noise, our performance is comparable to FG2, while under $\pm 180^{\circ}$ noise, our method outperforms CCVPE but falls short of SliceMatch.
@@
Second, we evaluate a practical scenario where only a relative depth model (Wang et al., 2022; Jiang et al., 2021; Wang & Liu, 2024) is available at inference. We plug in different relative depth models with our estimated correspondences, without retraining or finetuning. As shown in the bottom two rows of the known and unknown orientation entries in Tab. 2, relative depth increases localization error by less than 0.2 m. This flexibility in depth predictors highlights the practicality and robustness of our method for real-world deployment, especially when a compact depth predictor is required.

### v3

@@ Abstract
We propose an accurate and interpretable fine-grained cross-view localization method that estimates the 3 Degrees of Freedom (DoF) pose of a ground-level image by matching its local features with a reference aerial image. Unlike prior approaches that rely on global descriptors or bird’s-eye-view (BEV) transformations, our method directly learns ground–aerial image-plane correspondences using weak supervision from camera poses. The matched ground points are lifted into BEV space with monocular depth predictions, and scale-aware Procrustes alignment is then applied to estimate camera rotation, translation, and optionally the scale between relative depth and the aerial metric space. This formulation is lightweight, end-to-end trainable, and requires no pixel-level annotations. Experiments show state-of-the-art accuracy in challenging scenarios such as cross-area testing and unknown orientation. Furthermore, our method offers strong interpretability: correspondence quality directly reflects localization accuracy and enables outlier rejection via RANSAC, while overlaying the re-scaled ground layout on the aerial image provides an intuitive visual cue of localization performance.
@@ 1 Introduction
Recently, FG2 (Xia and Alahi, 2025) demonstrated the feasibility of establishing local feature correspondences between ground and aerial images by propagating BEV point matches to the image plane. However, performing the matching in BEV is suboptimal, since warping a ground image into BEV inevitably introduces ray-directional distortions (Song et al., 2024) and discards information along the height dimension (Xia and Alahi, 2025; Wang et al., 2024a). This loss of information hinders matching to aerial images and, as a result, degrades localization performance, especially in challenging settings such as unknown camera orientation.
Therefore, we propose to establish correspondences directly between ground and aerial images. Instead of relying on pixel-level annotations (Leroy et al., 2024; Wang et al., 2025), our approach is weakly supervised using only 3-DoF camera poses. To infer the pose from image-plane correspondences, we lift the sampled ground points into the BEV space with the help of monocular depth models (Yang et al., 2024a; b; Piccinelli et al., 2025; Wang and Liu, 2024). In unseen environments, usually only the non-metric relative depth prediction, i.e., depth up to an unknown scale, is reliable. Therefore, our method supports both metric and relative depth, and jointly estimates the camera pose and depth scale using scale-aware Procrustes alignment (Umeyama, 1991).
@@ 2 Related Work
Fine-grained cross-view localization is challenging due to the drastic viewpoint differences and asynchronous capture times between ground and aerial imagery. State-of-the-art approaches tackle this domain gap primarily by learning a deep network for camera pose estimation. Global descriptor-based methods (Xia et al., 2022; Lentsch et al., 2023; Xia et al., 2023) leverage contrastive learning to pull the ground and aerial features to a common representation space, and compare the ground image descriptor to aerial descriptors extracted at different candidate poses. Geometry transformation-based methods (Fervers et al., 2023b; Wang et al., 2023; Sarlin et al., 2024; Wang et al., 2024a; Song et al., 2024; Shi et al., 2023; Wang et al., 2024b; Fervers et al., 2023a; Xia and Alahi, 2025) first warp the ground image into a BEV representation, and then either perform a sliding-window search over the aerial image to find the best alignment (Sarlin et al., 2024; Fervers et al., 2023b; Shi et al., 2023) or estimate the relative pose by matching features between the BEV and aerial view (Song et al., 2024; Wang et al., 2024b; 2023; a; Xia and Alahi, 2025). Despite the increasing localization accuracy, most methods do not explicitly identify matched local features across views, offering limited interpretability. A recent work (Xia and Alahi, 2025) demonstrated, for the first time, the ability to find ground-aerial local feature correspondences by defining a dense 3D point cloud and querying image features for each point. However, this process is inefficient due to the sparse nature of 3D space, and this method performs poorly when camera orientation is unknown.
@@ 3.2 Camera Pose Estimation
Coordinate assignment: Given the $N$ sampled correspondences, we leverage off-the-shelf monocular depth estimators to assign coordinates to these correspondences, which are then used to estimate the camera pose. Importantly, monocular depth prediction is a geometrically ill-posed problem, as there is no unique mapping from a single image to absolute metric depth. State-of-the-art methods either learn generalizable metric depth from diverse training data (Piccinelli et al., 2025; Zhu et al., 2024), or predict depth that is accurate only up to an unknown per-image scale (Wang and Liu, 2024; Yang et al., 2024a). To generalize across diverse scenarios and image types, our method accommodates both metric and relative depth, and explicitly estimates a scale factor $s$ to convert relative depth into metric space. This scale estimation leverages the metric information, i.e., meters per pixel, available in geo-referenced aerial imagery.
@@
Scale-aware Procrustes alignment: As shown in Fig. 2 right, once we have the point correspondences $\{(x_{n}^{A},y_{n}^{A}),(x_{n}^{G},y_{n}^{G},z_{n}^{G})/s\}$, and their matching probabilities $w_{n}$, we can compute the rotation $\mathbf{R}$, translation $\mathbf{t}$, and the scale $s$ between the ground points and the metric space aerial points analytically in a differentiable manner using 2D scale-aware Procrustes alignment (Umeyama, 1991)  [^1] In Xia and Alahi (2025); Barroso-Laguna et al. (2024), all coordinates are in metric space, so they use orthogonal Procrustes (Gower, 1975) and ignore scale..
For simplicity, we denote the planar coordinates [^2] We assume that the ground camera’s viewing direction is orthogonal to the gravity direction, since this can be reliably calibrated (Veicht et al., 2024) or obtained from an IMU or accelerometer. of all ground points as $\mathbf{P}=\mathbf{P}^{*}/s$, assuming they differ from the metric coordinates, $\mathbf{P^{*}}=\{(x^{G}_{1},y^{G}_{1}),...,(x_{n}^{G},y_{n}^{G})\}$, by an unknown scale $s$. The coordinates of all aerial points are denoted as $\mathbf{Q}$. The objective is to estimate the scale $s$, rotation matrix $\mathbf{R}$ and metric translation $\mathbf{t}$ that satisfy the transformation $\mathbf{Q}=s(\mathbf{R}\cdot\mathbf{P})+{\mathbf{t}}$.
@@
Next, we perform Singular Value Decomposition (SVD) on $\mathbf{C}$, where $\mathbf{C}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$, and the optimal rotation can be recovered as $\mathbf{R}=\mathbf{V}\mathbf{U}^{\top}$, which derives the estimated yaw orientation $o$. The scale and translation from the ground points $\mathbf{P}$ to the aerial points $\mathbf{Q}$ are then computed as,
@@
Recall that $\mathbf{P}=\mathbf{P}^{*}/s$. Substituting this into Eq. 4, the estimated scale $s^{*}$ can be expressed using the (unscaled) metric coordinates $\mathbf{P}^{*}$, with $\mathbf{\Sigma}^{*}$ computed from the SVD between $\mathbf{Q}$ and $\mathbf{P}^{*}$:
@@ 3.3 Model Supervision
We follow Xia and Alahi (2025) to provide supervision on both the pose and correspondences. For camera pose, we adopt the Virtual Correspondence Error (VCE) loss. Specifically, we define a set of $N_{v}$ virtual points $\mathbf{P}_{v}$ in a 2D metric space, where the hyperparameter $l$ controls the spatial extent of the space. We apply both the ground-truth transformation, $\mathbf{R}_{\text{gt}},\mathbf{t}_{\text{gt}}$, and the estimated transformation, $\mathbf{R},\mathbf{t}$, to these virtual points. The loss $\mathcal{L}_{\text{VCE}}$ then minimizes the mean Euclidean distance between the corresponding transformed points,
@@ 4 Experiment
We first introduce the datasets and metrics, followed by implementation details. We then compare our method with prior state-of-the-art vision-based methods and showcase inference using different monocular depth predictors. Next, we present qualitative results on local feature matching and interpretable layout alignment, along with cross-dataset generalization and our ablation studies.
@@ 4.1 Datasets and Evaluation Metrics
KITTI (Geiger et al., 2013) provides forward-facing ground-level images with a limited field of view. Aerial images for KITTI were collected by Shi and Li (2022), who also split the dataset into Training, Test1, and Test2 subsets. Test1 includes images from the same region as the Training set, while Test2 contains images from a different unseen region. We refer to Test1 and Test2 as the same-area and cross-area test sets, respectively. We use the metric depth predictions from DepthAnythingV2 (Yang et al., 2024b) for our experiments on KITTI.
VIGOR (Zhu et al., 2021) is a widely used cross-view localization dataset, containing ground-level panoramas and aerial images from four U.S. cities. It also provides same-area and cross-area splits. In the same-area split, training and test images are collected from all four cities, while in the cross-area split, the training and test sets come from two different cities. During training, we adopt a recent metric depth model, Unik3D (Piccinelli et al., 2025), which has shown strong performance on outdoor panoramic images. At test time, we report results using both Unik3D and relative depth models (Wang and Liu, 2024; Jiang et al., 2021; Wang et al., 2022). Additionally, we include results in the Appendix where both training and testing are performed using only relative depth.
@@ 4.2 Implementation Details
On both datasets, we use AdamW (Loshchilov and Hutter, 2019) with a learning rate of $1\times 10^{-4}$. Training is conducted on a single H100 GPU with a batch size of 80 for VIGOR and 224 for KITTI. We follow the setting from Xia and Alahi (2025): the temperature parameter $\tau$ in matching score computation is set to $0.1$, the aerial feature map $\mathbf{F}_{A}$ is resized to generate $41\times 41$ aerial points, $\mathcal{L}_{\text{VCE}}$ uses $N_{v}=10\times 10$ points, with $l$ set to 5 m, and $N=1024$ correspondences are sampled for pose estimation. For $\mathcal{L}_{\text{S2G}}$, negatives are 1 m away in planar distance from the projected location. We use $\beta=1$ for VIGOR and $\beta=0.1$ for KITTI. RANSAC (Fischler and Bolles, 1981) is applied on VIGOR, but omitted on KITTI as it did not yield improvements in localization accuracy. When using metric depth, we apply a maximum depth threshold of 35 m for VIGOR and 40 m for KITTI, setting the matching score of points exceeding this threshold to zero. For relative depth predictors, we apply a fixed initial scale (identified by visually inspecting a few examples) to all relative depth maps and then use the threshold to filter out matches corresponding to sky and distant objects.
@@ 4.3 Quantitative Results
KITTI: We compare our method with both global descriptor-based methods, CCVPE (Xia et al., 2023) and SliceMatch (Lentsch et al., 2023), and geometry transformation-based methods, GGCVT (Shi et al., 2023), HC-Net (Wang et al., 2024b), DenseFlow (Song et al., 2024), and FG2 (Xia and Alahi, 2025). As shown in Tab. 1, our method sets a new state-of-the-art in mean and median localization errors on the cross-area test under both $\pm 10^{\circ}$ and $\pm 180^{\circ}$ orientation noise. On the same-area test, although our method has higher error under $\pm 10^{\circ}$ noise, it shows a large improvement over the previous state-of-the-art method in the more challenging $\pm 180^{\circ}$ setting, reducing the mean error from 6.88 m to 1.85 m. For orientation prediction, our method is less accurate than the state of the art. Since our method computes orientation analytically from local feature correspondences, it cannot take full advantage of the prior that most ground images in KITTI are aligned with the road direction. Under $\pm 10^{\circ}$ noise, our performance is comparable to FG2, while under $\pm 180^{\circ}$ noise, our method outperforms CCVPE but falls short of SliceMatch.
@@
Second, we evaluate a practical scenario where only a relative depth model (Wang et al., 2022; Jiang et al., 2021; Wang and Liu, 2024) is available at inference. We plug in different relative depth models with our estimated correspondences, without retraining or finetuning. As shown in the bottom two rows of the known and unknown orientation entries in Tab. 2, relative depth increases localization error by less than 0.2 m. This flexibility in depth predictors highlights the practicality and robustness of our method for real-world deployment, especially when a compact depth predictor is required.
@@ 4.4 Interpretability
Refer to caption(a)
Refer to caption(b)
Refer to caption(a)
Refer to caption(c)
Refer to caption(b)
Refer to caption(d)
Refer to caption(c)
Refer to caption(d)
Figure 6: Direct generalization to the CVACT dataset. We visualize the top 50 correspondences, ranked by matching score, and overlay the ground layout on the aerial image after using the predicted rotation, translation, and scale transformations. We use the metric depth prediction from Unik3D.
4.5 Cross-Dataset Generalization
Additionally, we demonstrate the strong generalization capability of our model by directly applying the model trained on VIGOR to the CVACT dataset (Liu and Li, 2019). CVACT is a cross-view image retrieval dataset collected in Canberra, Australia, where the rural and natural landscapes present a significant domain gap compared to VIGOR. Notably, CVACT lacks precise localization labels (Shi et al., 2022). Therefore, we provide local feature matching results and evaluate localization performance by overlaying the transformed ground layout onto the aerial image.
In general, our model exhibits strong generalization capability across datasets collected in different countries. As shown in Fig. 6, our model establishes reasonable feature correspondences. In samples (a) and (c), features extracted along trees in the ground view accurately match the corresponding blobs in the aerial view. In samples (b) and (d), the curbs and road markings in the ground view are correctly matched to those in the aerial view. Reliable feature matching leads to accurate localization. As shown, the transformed ground layout consistently aligns well with the aerial image across all samples. In (a) and (b), road markings and lane boundaries projected from the ground view form seamless continuations of those in the aerial view. In (c) and (d), both ground-level structures (e.g., curbs) and above-ground objects (e.g., buildings and trees) exhibit strong alignment.